# Study of Venera Spacecraft Trajectories and Wider Implications

### v1

@@ Abstract
Historically, there is no doubt that the early years of the USSR space program put them way ahead of the competition (the USA). Nonetheless, although this was not what the Russians wished to present to the world, the interplanetary campaign, centred around missions to the planet Venus (the Venera program) was also beset with difficulties. Many of the early Venera probes failed, despite making it to a heliocentric orbit, but naturally the success rate improved with time. The result is that there are now many Venera probes in heliocentric orbits, either completely intact, or the main bus after a successful deployment of the lander; together with the associated Blok-L upper stages. This paper is a response to some previous quite contentious research proposing that a certain member of a new class of objects, designated $2005\ VL_{1}$ may in fact be the Venera-2 probe. In this paper we look into the invariance of the Earth Tisserand parameter in an attempt to establish if there are indeed any members of this class which could be Venera probes. It is found, with extremely small probability, that compared to a sample of randomly chosen NEOs, members of the class of Dark Comet have an Earth Tisserand unusually close to 3, a property shared by the Venera missions. Furthermore there are particular associations of 3 Dark Comets with 3 of these probes, the most significant being $2010\ RF_{12}$ with the Venera-12 mission.
@@ 2.6 Comparison of Dark Comets with Venera probes
Now we refer to Figure 9 and we find that for $2005\ VL_{1}$, all of the Venera probes are displaced significantly from Tisserand of $2005\ VL_{1}$, and furthermore, so is the minimum deviation of the probes. Thus the question mark indicates that we cannot firmly attribute ANY of the Venera probes to this particular object. This is also the situation for $2010\ VL_{65}$ $2003\ RM$. For the remaining 3 Dark Comets, $2016\ NJ_{33}$, $2010\ RF_{12}$ and $2006\ RH_{120}$, the Veneras 6, 12, & 3 respectively are the most likely candidates, with deviations of 0.09%, 0.05% and 0.01% . All these, but especially the last 2, represent discrepancies in the Tisserands of around the alteration in its value expected from an Earth encounter, as determined empirically in Section 2.5 and shown in Figures 7 and 8.
@@
$2016\ NJ_{33}$ Venera-6 0.09% 1.5
$2010\ RF_{12}$ Venera-12 0.05% 0.75
$2006\ RH_{120}$ Venera-3 0.01% 0.17
@@ 3 Discussion
We find that the association of Venera-3 with Dark Comet $2006\ RH_{120}$ is not entirely unusual in that there are at least 9 other Veneras which could be similarly ascribed. For the association of Venera-6 to $2016\ NJ_{33}$ , this is just within the realms of possibility, since Venera-6 (and Venera-5 for that matter) are only marginally outside the $2\sigma$ limits determined in Section 2.5. Finally we can conclude that the association of $2010\ RF_{12}$ with Venera-12 is quite significant, and it is unlikely ANY other of the Veneras could be this object.
@@
Venera-6 launched on 1969 Jan 10th, with arrival on 1969 May 17th, having separated into a lander section and bus at a distance of $25,000$ $\mathrm{k}\mathrm{m}$ from Venus. The mission was a success it seems and the bus continued on into a heliocentric orbit. The value of $R_{NUC}$ for $2016\ NJ_{33}$ in Seligman et al. (2023) is given as $\sim{16}$ $\mathrm{m}$, though the caveat mentioned for $2010\ RF_{12}$ also applies here.
Finally Venera-3. Launched on 1965 Nov 16th, and impacted Venus on 1966 Mar 1st after a course correction on 1965 Dec 26th. However as already articulated, this does not preclude $2006\ RH_{120}$ from being associated with Venera-3 in some way, since the Blok-L for this mission may be an alternative candidate. The value of $R_{NUC}$ for $2006\ RH_{120}$ is in the range 2-7 $\mathrm{m}$. Now let us examine the three Dark Comets in more detail.

### v2

@@ Abstract
Historically, there is no doubt that the early years of the USSR space program put them way ahead of the competition (the USA). Nonetheless, although this was not what the Russians wished to present to the world, the interplanetary campaign, centred around missions to the planet Venus (the Venera program) was also beset with difficulties. Many of the early Venera probes failed, despite making it to a heliocentric orbit, but naturally the success rate improved with time. The result is that there are now many Venera probes in heliocentric orbits, either completely intact, or the main bus after a successful deployment of the lander; together with the associated Blok-L upper stages. This paper is a response to some previous quite contentious research proposing that a certain member of a new class of objects, designated $2005\ VL_{1}$ may in fact be the Venera-2 probe. In this paper we look into the invariance of the Earth Tisserand parameter in an attempt to establish if there are indeed any members of this class which could be Venera probes. It is found, with extremely small probability, that compared to a sample of randomly chosen NEOs, members of the class of Dark Comet have an Earth Tisserand unusually close to 3, a property shared by the Venera missions. Furthermore there are particular associations of 3 Dark Comets with 3 of these probes, the most significant being $2010\ VL_{65}$ with the Venera-12 mission.
@@ 2.6 Comparison of Dark Comets with Venera probes
Now we refer to Figure 9 and we find that for $2005\ VL_{1}$, all of the Venera probes are displaced significantly from Tisserand of $2005\ VL_{1}$, and furthermore, so is the minimum deviation of the probes. Thus the question mark indicates that we cannot firmly attribute ANY of the Venera probes to this particular object. This is also the situation for $2010\ RF_{12}$ $2003\ RM$. For the remaining 3 Dark Comets, $2016\ NJ_{33}$, $2010\ VL_{65}$ and $2006\ RH_{120}$, the Veneras 3, 12, & 6 respectively are the most likely candidates, with deviations of 0.014%, 0.05% and 0.05% . All these, but especially the last 2, represent discrepancies in the Tisserands of around the alteration in its value expected from an Earth encounter, as determined empirically in Section 2.5 and shown in Figures 7 and 8.
@@
$2016\ NJ_{33}$ Venera-3 0.014% 0.23
$2006\ RH_{120}$ Venera-6 0.05% 0.75
$2010\ VL_{65}$ Venera-12 0.05% 0.75
@@ 3 Discussion
We find that the association of Venera-3 with Dark Comet $2016\ NJ_{33}$ is not entirely unusual in that there are at least 6 other Veneras which could be similarly ascribed. For the association of Venera-6 to $2016\ RH_{120}$ , this is certainly within the realms of possibility, since Venera-6 (and Venera-5 for that matter) are just inside the $2\sigma$ limits determined in Section 2.5. Finally we can conclude that the association of $2010\ VL_{65}$ with Venera-12 is quite significant, and it is unlikely ANY other of the Veneras could be this object.
@@
Venera-6 launched on 1969 Jan 10th, with arrival on 1969 May 17th, having separated into a lander section and bus at a distance of $25,000$ $\mathrm{k}\mathrm{m}$ from Venus. The mission was a success it seems and the bus continued on into a heliocentric orbit. The value of $R_{NUC}$ for $2016\ RH_{120}$ in Seligman et al. (2023) is in the range 2-7 $\mathrm{m}$, though the caveat mentioned for $2010\ VL_{65}$ also applies here.
Finally Venera-3. Launched on 1965 Nov 16th, and impacted Venus on 1966 Mar 1st after a course correction on 1965 Dec 26th. However as already articulated, this does not preclude $2016\ NJ_{33}$ from being associated with Venera-3 in some way, since the Blok-L for this mission may be an alternative candidate. The value of $R_{NUC}$ for $2016\ NJ_{33}$ is $\sim{16}$ $\mathrm{m}$. Now let us examine the three Dark Comets in more detail.